# Communicating the gravitational-wave discoveries of the LIGO–Virgo–KAGRA Collaboration

### v2

@@ 2.7 Multisensory resources: sound and touch
GWs can also be communicated through touch. The tidal spatial deformations of the first detected GW signal have been translated to a 3D-printable model, specifically tailored to the needs of visually impaired people. By running fingers along the edges of the shape, one can appreciate the features of a GW and perceive the details of the signal evolving in time. This design is freely available [21], along with suggestions for classroom activities.
@@
Figure 4: A Tactile Universe 3D printed surface of the inspiral of two black holes and the resulting GWs (credit: Coleman Krawczyk).
@@
The Tactile Universe (www.tactileuniverse.org[22] has partnered with UK LVK groups to design tactile resource sets for use with blind and vision-impaired pupils in upper secondary education. A mix of 3D-printable (Figure 4) and more basic tactile resources, paired with existing sonifications, are tied to 20 minute workshops that teach students about black holes and neutron stars, GW signals and detection, and core science topics like waves and gravity. The 3D-printable resources, example lesson scripts and guides are available on the website.
2.8 Art and music with science
Numerous art, music, and science collaborations explore novel approaches and interpretations of GWs. These collaborations showcase our science through a different and interesting point of view and often bring the topic of GWs to new audiences. Works have been created by both professional artists and musicians and members of the LVK [23].
Just a few examples include music compositions [24, 25], open land art installations [26], museum exhibits [27] and novel musical instruments (gravitysynth.tumblr.com). The LIGO-India blog has a virtual gallery GWSciArt that showcases GW-inspired artistic projects [28].
2.9 Social media and non-traditional media
@@
Non-traditional media provides an opportunity to present our science to broad audiences. Activities like social-media posts, blogs, and podcasts or interviews can help communicate science in an informal way, e.g., the Spanish podcasts “Oscilador armónico”  [29], “A ciencia cierta” [30] and “Coffee Break: Señal y Ruido” [31] which coined and popularized the term “gravitondas” (“gravit-waves”) for GWs.
The LVK has several social media accounts, for the collaborations and the observatory sites, on X, Facebook, Instagram, Mastodon (the Astrodon instance). Content includes educational resources, news stories, and discovery announcements. Special posts are scheduled for events such as detection anniversaries or the International Day of Women and Girls in Science. Reddit has been used for Ask-me-anything posts about discoveries. Social media allows the LVK to interact directly with the public, and threads in response to questions have proved popular.
2.10 Computer games/apps, visualisations, and interactive technology
@@
Interactive apps and games are engaging and fun ways to convey GW science. Physical interaction with virtual environments incorporates additional modes for audiences to engage with content. Activities range from computer games [32] (www.laserlabs.org/; blackholehunter.org) to classroom and science festival teaching with virtual reality [33, 34] (www.scivr.com.au).
@@ 3 Case Study: GWTC-3
GW discoveries from the LVK are published in catalogs: papers accompanied by data releases. GWTC-3 [3] is the most recent of these, presenting results up to the end of our third observing run. Thanks to the continued improvements in detector sensitivity [35], the rate of discoveries has increased with time, meaning that each catalog includes a significant number of new GW candidates. GWTC-3 represents the most sophisticated and comprehensive analysis published by the LVK to date.
Production and publication of a catalog is a multi-year project directly involving hundreds of experts. Given the significance of the results and the importance of easy-to-use data releases, effective communication is a priority. However, communication of catalog results faces: catalog results consist of comprehensive analyses of many signals (making them overwhelming), and the majority of detections have properties similar to those observed in the past (lacking the excitement of a novel discovery). For GWTC-3, the project team included a member responsible for coordinating communication and outreach activities from the beginning. This embedding allowed for these activities to be planned as the project progressed, developed in tandem with the data releases, and coordinated with those compiling the results.
@@
3.1 Academic
A series of LVK webinars were organised to coincide with the release of GWTC-3. The first on GWTC-2.1 [36], a reanalysis of the previous data using updated methods consistent with GWTC-3. Subsequent webinars in the series presented the GWTC-3 results and implications for cosmology, astrophysics and tests of general relativity. Presenters were drawn from those who made key contributions to the work, with preference given to early-career scientists, and ensuring a balance in geographic location and demographics.
The data release was a key component of the GWTC-3 results. It included Jupyter notebooks and Python scripts to demonstrate use and how to reproduce plots from the paper. Additionally, a Streamlit app was created to interactively make plots of source parameters: this enabled users to customise paper plots without needing to download data (gwtc3-contours.streamlit.app/).
3.2 Written resources
Like the paper itself, the GWTC-3 Science Summary [37] had a broad scope to cover. It was written in close collaboration between outreach experts and the analysis and paper-writing teams. It includes background information needed to understand the catalog, and a subset of highlight discoveries. The summary has been translated into $9$ languages.
3.3 Magazine article
@@
LIGO Magazine issue $20$ featured an $8$ page article on the story behind the catalog [38]. Short snippets from $23$ members of the analyses, paper writing, and engagement teams wrote about their experiences and stories of working on the catalog and associated results. Although many individuals contributed to this feature, they represent only a fraction of those providing input to GWTC-3. Articles like this give readers a behind-the-scenes insight of working on big LVK results.
3.4 Graphics and visualisations
GWTC-3 updates to some of our long-standing visualisations were made. The new results were added to the Masses in the Stellar Graveyard plot  5), showing the masses of black holes and neutron stars observed by the LVK alongside those observed via traditional electromagnetic astronomy. An online version allows for customisable images [39]. The LVK Orrery was also updated, which shows a stylised animation of the GWTC-3 binaries.
New visualisations were also created. One was a poster reminiscent of a periodic table showing the catalog after each observing period. The poster was used in a variety of articles, a double-page LIGO Magazine spread [40], and in the GWTC-3 webinar. An artistic representation of the detections in a box accompanied the publication in Physical Review X. An image showing spectrograms for the 90 GWTC-3 signals was featured as NASA Astronomy Picture of the Day (7 December 2021).
@@ 3.4 Graphics and visualisations
Figure 5: GWTC-3 version of Masses in the Stellar Graveyard (credit: LVK/A Geller/Northwestern). The plot shows the known masses of stellar-mass black holes and neutron stars observed electromagnetically and with GWs.
3.5 Audification
@@
Low Mass Beats is a audification of the catalog, converting masses into pitch for each black hole or neutron star: the greater the mass, the lower the pitch. This lighthearted project provides an entertaining perspective on the catalog (SoundCloud: soundcloud.com/user-889003031).
3.6 Social media
@@
GWTC-3 material was shared via a social media campaign in two bursts using the hashtag #GWCT3. The first burst coincided with the release of the preprint and associated data. Special attention was given to advertising the webinars and data releases. The second burst coincided with the publication of the journal article and priority was given to any resources developed since the initial release, such as the Streamlit plotter and Low Mass Beats.
@@
For both bursts, posts were spaced over multiple days to catch the attention of people in different timezones. Material was translated into regional languages [41]. The pacing also made it natural to integrate popular-science articles written immediately after the release.
@@
Resources were subsequently reshared, e.g., Low Mass Beats was shared on Halloween because of its spooky sound, and the Orrery was shared during NASA’s Black Hole Week.
@@
The impact of the social media campaign can be partially assessed via the Altmetric score. As of June 2024, the GWTC-3 paper has a score of $713$ (including $56$ news outlets and 5 Wikipedia pages). This score is in the $99$th percentile of all tracked outputs.
@@ 4 Conclusions
LVK activities span a range of audiences and media. Many resources are used in multiple contexts, such that utility can be amplified beyond the original target, e.g., paper plots used on social media, infographics in talks, and data audifications in musical performances.
@@
  • GWs are an unfamiliar means of performing astronomy. Unlike traditional astronomy, we do not directly image our sources. However, GW data is suitable for audification, and a variety of tactile resources have been made. These make GW astronomy accessible to people with vision impairments, who are often excluded from accessible astronomy communication [42].
  • LVK results are produced by large teams. This makes it difficult for those outside to understand how discoveries are made, or to construct human-interest stories. However, this enables us to emphasise the importance of teamwork and international cooperation to science. The LVK’s breakthroughs are a direct contrast to the popular perception that scientific breakthroughs come from lone geniuses [43, 44]. The diversity of the LVK’s membership allows us to draw upon many different lived experiences, and demonstrate that scientists may come from any background.
  • After the initial first observations, many detections were similar to those discovered previously. The increase in the LVK’s detection rate is an experimental triumph—what was once extraordinary is now everyday. However, many detections are now not novel. It is therefore necessary to concentrate on the science enabled by a large set of detections. This has a benefit of reflecting that most scientific progress comes from careful study rather than a dramatic breakthrough.
While the LVK has benefitted from an existing interest in black holes and space, it has also faced hurdles, such as a tendency to misattribute discoveries to LIGO alone. As the field continues to mature, we expect that communication strategies will need to continue to evolve to reflect both the state of the field, and the interest and awareness of our audiences.

### v3

@@ 2.7 Employing multisensory resources
GWs can also be communicated through touch. The GW150914 signal has been translated to a 3D-printable model, specifically tailored to the needs of visually impaired people. By running fingers along the edges of the shape, one can appreciate the features of a GW and perceive the details of the signal evolving in time. This design is freely available [EGO-Virgo, 2021], along with suggestions for classroom activities.
@@
Figure 5: A Tactile Universe 3D printed surface of the inspiral of two black holes and the resulting GWs
@@
The Tactile Universe (www.tactileuniverse.org[Bonne et al., 2018] has partnered with LVK groups to design tactile resource sets for use with blind and vision-impaired pupils in upper secondary education. A mix of 3D-printable (Figure 5) and more basic tactile resources, paired with existing sonifications, are tied to $20$ minute workshops about black holes and neutron stars, GW signals and detection, and core science topics like waves and gravity. The 3D-printable resources, example lesson scripts and guides are available on the website. Students have enjoyed interacting with the tactile resources; teachers have praised the accessibility and pacing of the workshops, and in post-session feedback, students have demonstrated an understanding of the concepts discussed.
The audifications and tactile resources have also proved effective at science festivals, with members of the public appreciating having the diverse means of explaining unfamiliar concepts. This demonstrates how providing multisensory resources makes ideas accessible to a wider audience as well as more comprehensible to existing audiences.
2.8 Combining art and music with science
Art, music, and science collaborations explore novel approaches to GW communication, showcase the science through a different point of view, and bring the topic to new audiences. Works have been created by professional artists and musicians, and LVK members [e.g., Azure et al., 2021]. Examples include music compositions [Instituto de Física Corpuscular, 2023; Penguin Cafe, 2023], open land art installations [Virgo, 2022], museum exhibits [EGO, 2019] and novel musical instruments such as GravitySynth which combines the technology of GW detection with modular synthesisers [Trimble, 2024; Azure et al., 2021]. The LIGO-India blog has a virtual gallery GWSciArt that showcases GW-inspired artistic projects [LIGO-India, 2024c].
Combining art with science can help reach new audiences and provide a memorable experience. For example, the art and science festival Celebrating Einstein merged dance, music, and film with GWs. Assessment of the festival audience demonstrated that those who attended were a mix of those interested in science and art; that non-physics experts gained knowledge (e.g., $75\%$ of physics novices improved their score between pre- and post-event surveys), and that attendees typically had a positive emotional response to the event [Grimberg et al., 2019]. Similarly, Into the Quadrivium, a collaboration between GW researchers and musicians, which blended baroque and contemporary music with spoken-word explanations of GW science, was positively received by the audience [Into the Quadrivium, 2023].
2.9 Communicating through social and non-traditional media
@@
Non-traditional media provides an opportunity to present our science to broad audiences. Activities like social-media posts, blogs, and podcasts or interviews can communicate science in an informal way.
Interviews have been hosted on the Spanish podcasts “Oscilador armónico”  [Cordero-Carrión et al., 2023] ($7,800$ listens as of August 2024), “A ciencia cierta” [Rivera, 2021] ($33,100$ listens) and “Coffee Break: Señal y Ruido” [Socas-Navarro, 2021] ($40,800$ listens). The last podcast coined the term “gravitondas” (“gravi-waves”) for GWs. Communicating with audiences in their local language makes science more accessible. The podcast format is more flexible than traditional radio programming allowing more time and in-depth discussion needed to explain complicated ideas.
The LVK has several social-media accounts, for the Collaborations and the observatory sites, across different platforms (Table 2). The greater following for LIGO compared to Virgo and KAGRA may be a consequence of the same bias that leads to LIGO preferentially receiving credit. Content includes educational resources, news stories, and discovery announcements. Special posts are scheduled for events such as detection anniversaries or the International Day of Women and Girls in Science. The observatory accounts often share posts about local news, such as pictures from the sites or events at the science centres, while the Collaboration accounts take the lead on big announcements.
Table 2: Social-media followers (August 2024).
Platform Account Followers
X LIGO Scientific Collaboration $110,600$
European Gravitational Observatory and Virgo $12,500$
KAGRA $2,500$
LIGO Hanford Observatory $10,800$
LIGO Livingston Observatory $5,300$
LIGO India $6,700$
LIGO Magazine $690$
Facebook LIGO Scientific Collaboration $33,000$
European Gravitational Observatory and Virgo $7,000$
LIGO Hanford Observatory $6,100$
LIGO Livingston Observatory $6,800$
LIGO India $9,600$
Instagram LIGO–Virgo $14,900$
LIGO India $5,600$
Mastodon (Astrodon) LIGO Scientific Collaboration $1,500$
Social-media content is reshared across platforms with suitable adaptations. The microblogging X and Mastodon platforms are well-suited to threads; these are useful for linking many resources or explaining a topic from many angles. Facebook and Instagram allow for longer posts, enabling more in-depth explanation per post, but posts must be spaced further apart to avoid cluttering followers’ feeds. Across all platforms, the most popular posts often feature a graphic or video (all Instagram posts include a graphic or video), and hence these are sought whenever possible.
The different social-media platforms reach different numbers of people and demographics. Table 3 shows metrics for the GW230529_181500 discovery announcement (April 2024). While X has the most views/users reached; Mastodon has the most shares per follower (as a newer platform it potentially has a higher proportion of active users), and Instagram has the most likes per follower.
Table 3: Social-media performance as of August 2024 for LIGO posts about GW230529_181500 (GW230529) and the GWTC-3 Orrery. Statistics are not directly comparable between platforms but qualitatively similar. Numbers per follower are given in parentheses.
Platform
Post Metric X Facebook Instagram Mastodon
GW230529 Reposts/shares/boosts $155$ ($0.0014$) $89$ ($0.0027$) $72$ ($0.0048$) $18$ ($0.0120$)
Likes/reactions/favorites $328$ ($0.0030$) $261$ ($0.0079$) $662$ ($0.0444$) $26$ ($0.0173$)
Views/reach $46,000$ ($0.42$) $18,300$ ($0.55$) $5,200$ ($0.35$) n/a
Orrery Reposts/shares/boosts $113$ ($0.0010$) $48$ ($0.0014$) $51$ ($0.0034$) $25$ ($0.0167$)
Likes/reactions/favorites $350$ ($0.0032$) $144$ ($0.0043$) $813$ ($0.0546$) $32$ ($0.0213$)
Views/reach $192,000$ ($1.74$) $7,800$ ($0.24$) $9,600$ ($0.64$) n/a
Social media allows the LVK to interact directly with the public, and threads in response to questions have often proved popular. On Reddit, the LVK has run occasional Ask-me-anything discussions. The Ask-me-anything accompanying GW150914 has $2,300$ comments. Posting answers to questions online enables them to be searched out in the future.
2.10 Using interactive technologies
@@
Interactive apps and games are engaging and fun ways to convey GW science. Interactive apps and games are engaging and fun ways to convey GW science. Physical interaction with virtual environments incorporates additional modes for audiences to engage with content. Activities have ranged from computer games [Carbone et al., 2012] (www.laserlabs.org; blackholehunter.org) to classroom and science festival teaching with virtual reality (www.scivr.com.au[Parks, 2023] .
Mission Gravity [Bondell & Myers, 2021; Kersting et al., 2023] is an interactive virtual-reality environment designed for secondary school students to collaborate to better understand black holes and neutron stars. Since 2018, the program has been delivered to over 23,000 students across over $400$ schools, as well as to over 1,100 teachers through professional-development workshops. One student reported: “The virtual reality component of it was engaging and enjoyable […] I was able to understand ideas that I previously struggled with.”
@@ 3 Case Study: GWTC-3
GW discoveries from the LVK are published in catalogs: papers accompanied by data releases. GWTC-3 [Abbott et al., 2023a] is the most recent of these, presenting results up to the end of the third observing run. Thanks to the continued improvements in detector sensitivity [Abbott et al., 2020b], the rate of discoveries has increased with time, meaning that each catalog includes a significant number of new GW candidates. GWTC-3 represents the most sophisticated and comprehensive analysis published by the LVK to date.
Production and publication of a catalog is a multi-year project directly involving hundreds of experts. Given the significance of the results and the importance of easy-to-use data releases, effective communication is a priority. However, communication of catalog results faces challenges: catalog results consist of comprehensive analyses of many signals (making them overwhelming), and the majority of detections have properties similar to those observed in the past (lacking the excitement of a novel discovery).
For GWTC-3, the project team included a member responsible for coordinating communication and outreach activities from the beginning. This embedding allowed for these activities to be planned as the project progressed, developed in tandem with the data releases, and coordinated with those compiling the results. Delaying this work until the project was nearly complete may not have left sufficient time to produce and review resources, and risked the team being too exhausted to contribute to these tasks. LVK paper teams typically have a designated person responsible for producing data products and for writing the Science Summary, but these may not necessarily be integrated into the team throughout the project, and there is not typically one person responsible for coordinating communication and outreach more generally. We recommend following an approach similar to GWTC-3 of having a dedicated team member responsible for communication and outreach, and starting work on the resources early.
@@
3.1 Sharing the people behind the science
@@
LIGO Magazine issue $20$ featured an $8$ page article by $23$ members of the analyses, paper-writing, and engagement teams [Berry et al., 2020]. They wrote about their experiences of working on the catalog and associated results. Although the article authors represent only a fraction of those providing input to GWTC-3, features like this give readers a behind-the-scenes insight of working on big LVK results.
3.2 Engaging academia
A series of LVK webinars were organised to coincide with the release of GWTC-3. The first on GWTC-2.1 [Abbott et al., 2024], a reanalysis of previous data using updated methods consistent with GWTC-3. Subsequent webinars presented the GWTC-3 results and implications for cosmology, astrophysics and tests of general relativity. Presenters were drawn from those who made key contributions to the work, with preference given to early-career scientists, and balancing geographic location and demographics.
The data release was a key component of the GWTC-3 results. It included Jupyter notebooks and Python scripts to demonstrate use and how to reproduce plots from the paper. As of August 2024, the data release for the inferred source properties of the new GWTC-3 detections is the most viewed of all Zenodo-hosted LVK data releases, with approximately $8,000$ views ($27,500$ file downloads); the second most viewed is the equivalent set of results for the GWTC-2.1 analysis, with approximately $5,600$ views ($21,800$ file downloads). The popularity of the data release indicates that the community appreciates the value of the results, and understands how to use data products.
As the detection rate increases, it becomes more difficult to manage all the analyses, and curate data releases that include all the relevant results (especially if there have been multiple reanalyses to obtain final results). We recommend automating the process to reduce the chance of human error in data-release production, and efforts are underway to develop such automation tools [e.g., Williams et al., 2023].
For GWTC-3, a Streamlit app was created to interactively make plots of source parameters enabling users to customise paper plots without needing to download data (gwtc3-contours.streamlit.app/).
3.3 Writing reference texts
Like the paper itself, the GWTC-3 Science Summary [LIGO-Virgo-KAGRA, 2021] had a broad scope to cover. It was written in close collaboration between outreach experts and the analysis and paper-writing teams. It includes background information needed to understand the catalog, and a subset of highlight discoveries. The summary has been translated into $9$ languages.
A LVK news item, rather than a press release, was drafted for GWTC-3. This was used by individual institutions to draft their press releases. The decision to not have a LVK press release reflected the expectation that GWTC-3 would be more difficult to communicate, and hence less attractive to write about, than a single discovery. Despite this, several science journalists did write articles about GWTC-3 [e.g., Carlise, 2021; Castelvecchi, 2021; Plait, 2021]. This demonstrates that the LVK did at least partially succeed in efforts to communicate the science of GWTC-3 to journalists, and (assuming similar communications efforts are organised in the future) that there may be an audience for press releases for future catalog releases.
Resources, including the news item and institutional press releases, were made public at 01:00 UTC on Monday 8 November 2023. The time was chosen to coincide with the paper preprint appearing on arXiv, while the date was chosen to correspond to the predefined detector data release timeline. An embargo period was offered from 18:00 UTC on Thursday 4 November when trusted journalists could be contacted about the upcoming release. Providing such an embargo period gives time for journalists to prepare and check their stories, which is beneficial; however, this embargo period was mostly over the weekend. Therefore, for future discovery announcements, it may be beneficial to have a longer embargo period, or to choose the day of the week for release more carefully.
3.4 Creating graphics
@@
Figure 6: GWTC-3 version of Masses in the Stellar Graveyard (credit: LVK/Aaron Geller/Northwestern). The plot shows the known masses of stellar-mass black holes and neutron stars observed electromagnetically and with GWs.
Updates to some long-standing LVK visualisations were made for GWTC-3. The new results were added to the Masses in the Stellar Graveyard plot (Figure 6). This plot is now generated from GWOSC data products, making it easy to add large numbers of detections. An online version allows for customisable images [Geller et al., 2024]. The LVK Orrery was also updated, which shows a stylised animation of the GWTC-3 binaries.
New visualisations were also created. One was a poster reminiscent of a periodic table showing the catalog after each observing period. The poster was used in a variety of articles, a double-page LIGO Magazine spread [Knox, 2022], and in the GWTC-3 webinar. An artistic representation of the detected sources in a box accompanied the publication in Physical Review X (Figure 7). An image showing spectrograms for the $90$ GWTC-3 signals was featured as NASA Astronomy Picture of the Day (7 December 2021).
Figure 7: An artistic representation of GWTC-3 sources which accompanied the journal publication (credit: Carl Knox/OzGrav Swinburne).
3.5 Employing multisensory resources
@@
Low Mass Beats is a audification of the catalog, converting masses into pitch for each black hole or neutron star: the greater the mass, the lower the pitch. This lighthearted project provides an entertaining perspective on the catalog. (SoundCloud: soundcloud.com/user-889003031)
3.6 Communicating through social and non-traditional media
@@
GWTC-3 material was shared via a social-media campaign in two bursts using the hashtag #GWTC3. The first burst coincided with the release of the preprint and associated data. Special attention was given to advertising the webinars and data releases. The second burst coincided with the publication of the journal article and priority was given to any resources developed since the initial release, such as the Streamlit plotter and Low Mass Beats.
@@
For both bursts, posts were spaced over multiple days to catch the attention of people in different timezones. Material was translated into regional languages [LIGO-India, 2024d]. The pacing also made it natural to integrate popular-science articles written immediately after the release.
From the second burst, the most popular post was for the Orrery. Statistics are shown in Table 3. It is generally less shared than the GW230529_181500 post, but there were more posts for GWTC-3 than for GW230529_181500 spreading out attention: that numbers are comparable demonstrates that catalog results can capture interest similarly to new detections. The Orrery has high X views having been reposted by accounts with large followings.
@@
Resources were subsequently reused, e.g., Low Mass Beats was shared on Halloween because of its spooky sound, and the Orrery was shared during NASA’s Black Hole Week.
@@
The impact of the social-media campaign can be partially assessed via the Altmetric score. As of August 2024, the GWTC-3 paper has a score of $714$ (including $56$ news outlets and $6$ Wikipedia pages). This score is in the $99\rm{th}$ percentile of all tracked outputs.
@@ 4 Conclusions
LVK activities span a range of audiences and media. Within the LVK, different activities are run by different teams, some coordinated Collaboration-wide, while others reflect individual interests. This provides a rich environment, where science-communication resources can be shared between different communications efforts. Many resources are used in multiple contexts, such that utility can be amplified beyond the original target, e.g., paper plots used on social media, infographics in talks, and data audifications in musical performances.
@@
  • GWs are an unfamiliar means of performing astronomy. Unlike traditional astronomy, GW astronomy does not directly image sources. However, GW data is suitable for audification, and a variety of tactile resources have been made. These make GW astronomy accessible to people with vision impairments, who are often excluded from accessible astronomy communication [Bell & Silverman, 2019], in addition to providing a novel way to explain these physical concepts.
  • LVK results are produced by large teams. This makes it difficult for those outside to understand how discoveries are made, or to construct human-interest stories. However, this enables emphasization of the importance of teamwork and international cooperation to science. The LVK’s breakthroughs are a contrast to the popular perception that scientific breakthroughs come from lone geniuses [Larivière et al., 2015; Aksnes & Aagaard, 2021]. The diversity of the LVK’s membership also allows us to draw upon many different lived experiences, and demonstrate that scientists may come from any background.
  • After the initial first observations, many detections were similar to those discovered previously. The increase in the LVK’s detection rate is an experimental triumph—what was once extraordinary is now everyday. It is therefore necessary to concentrate on the science enabled by a large set of detections. This has a benefit of reflecting that most scientific progress comes from careful study rather than a dramatic breakthrough.
Elements of each challenge may translate to other communication efforts:
  • When abstract or theoretical concepts are being communicated, having diverse explanations (combining multiple views of the concept and engaging multiple senses) may aid audience comprehension.
  • When results come from large collaborations, which is increasingly common in physics and astronomy [Smith, 2016; Battiston et al., 2019], drawing upon the diverse experiences of collaboration members may help engage members of the public with similar backgrounds.
  • When a new field is founded, charting the way the field will evolve long term may help increase scientific literacy regarding how scientific progress is made through many small advancements (in addition to through big discoveries).
Consequently, while we do not expect the exact circumstances faced by the LVK will be replicated, we anticipate that insights into GW-science communication may benefit others across science.
In communicating GW discoveries, the LVK has benefitted from existing interest in black holes and space. It has also faced hurdles, such as a tendency to misattribute discoveries to LIGO alone. This highlights the importance of taking a holistic approach to planning and reviewing communication: it is advantageous that an audience correctly understands how science is done, as well as what the discoveries are, and it cannot be assumed that just because the audience is interested they will absorb all relevant information.
As GW astronomy continues to mature, we expect that communication strategies will need to evolve to reflect both the state of the field, and audience interest and awareness.